THE SYMPOSIUM: A PERSONAL PERSPECTIVE

\underline{ROBERT CURL}, Department of Chemistry and Rice Quantum Institute, Rice University, Houston, TX 77005, USA.

The International Symposium on Molecular Spectroscopy at Ohio State University has been serving the spectroscopic community magnificently for sixty-five years. Many spectroscopists, myself included, made their very first meeting presentation here. In my case, that was a mere fifty-two years ago. Looking back my trips here have been filled with great experiences both personally and professionally. I plan to talk about a few of them.